OpenStudy (anonymous):

if you have integral of 1/(2x+3y)^5 dA R=[0.1]*[1,2] how do we solve it

OpenStudy (experimentx):

R=[0.1]*[1,2] ??

OpenStudy (experimentx):

this must be a definite integral ... what are the limits for x and y??

OpenStudy (anonymous):

iterated integral limits for x are [0,1] and y are [1,2]

OpenStudy (experimentx):

change to polar coordinates

OpenStudy (anonymous):

how

OpenStudy (experimentx):

\( x = r \cos \theta \; \; and \; \; y = r \sin \theta \)

OpenStudy (experimentx):

and get rid to those 2's and 3's

OpenStudy (experimentx):

and plot those points .... on your graph paper .. get limits for r and \theta

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \int\limits_{1}^{2} 1/(2x+3y)^2 dydx\] this is the equation but what happens when the power is changed

OpenStudy (experimentx):

Jeez ... you were making things complicated on your own??? I don't know ... for 2 i guess you will get the answer in same pattern ... physically ... they are whole lotta different ... can't say anything about the answer without doing it first

OpenStudy (anonymous):

how do i go abt doing it

OpenStudy (experimentx):

put x and y as i mentioned above .... and dA = r dr d\theta though i doubt if you can do in using rectangular coordinates

OpenStudy (anonymous):

i took the (2x+3y)^2 up to give me (2x+3y)^-2 whats the next step

OpenStudy (experimentx):

not sure about rectangular coordinates ... this is how i do it in polar coordinates \( 2x +3y = 2r\cos \theta + 3 r \sin\theta = r(2\cos\theta + 3\sin\theta) = r A\cos(\theta - \phi)\) where, \( A = \sqrt{2^2 + 3^3} \) and \( \phi = \tan^{-1} (2/3) \)

OpenStudy (anonymous):

im not allowed to use polar coordinates we have not reached that stage apparently its rectangular coordinates

OpenStudy (experimentx):

I thought you would do polar coordinate on High school and double integrals on undergraduate ... well ... let's see

OpenStudy (experimentx):

first integrate x or y treating other as constant!!

OpenStudy (anonymous):

final answer should be [-1/6 (\ln 8-\ln 2-\ln 5)\]

OpenStudy (experimentx):

Looks like pretty simple ... http://www.wolframalpha.com/input/?i=integrate+%282x+%2B+3b%29^-2+dx I was the one making things unnecessarily complicated

OpenStudy (anonymous):

will check it out its still loading

OpenStudy (experimentx):

evaluate this first \[ \int_0^1 \frac{1}{(3x+2y)^2} dx\]

OpenStudy (experimentx):

let \( 2x+3y = u, du = 3dx\) then we have \( u = 3y \) to \( 3y + 4\) \[ \frac{1}{3}\int_{3y}^{3y+4} \frac{du}{u^2}\]

OpenStudy (experimentx):

Oops ... sorry about the that ... 1/3 was supposed to be 1/2 ... I made mistake in copying the integral ... in the post two steps above

OpenStudy (anonymous):

im lost let me redo it ur way

OpenStudy (anonymous):

its 2x+3y

OpenStudy (anonymous):

k thnx

OpenStudy (anonymous):

y is ur du=3dx

OpenStudy (experimentx):

because i supposed 2x+3y=u, 2dx=du <--- it was supposed to be ... i made error copying!!

OpenStudy (anonymous):

ok where is the second u coming from?

OpenStudy (experimentx):

?? can you use draw function to demonstrate what you mean or ... clarify what you are trying to say??

OpenStudy (anonymous):

let 2x+3y=u,du=3dx then we have u=3y to 3y+4 that u=3y where do u get it bcz when i substitute back i get \[1/2 \int\limits_{0}^{1} 1/u^2 du\]

OpenStudy (anonymous):

what do i do after that step whats the integral of 1/u^2

OpenStudy (experimentx):

\[ \left [ \frac{u^{-1}}{-1}\right ]_{3y}^{3y+4} = \frac{1}{3y} - \frac{1}{3y+4}\] Integrate it again with respect to y \[ \int_1^2 \left ( \frac{1}{3y} - \frac{1}{3y+4} \right )dy\]

OpenStudy (experimentx):

Integrate 1/u^2 within those limits ... and get the result in y, and integrate it with dy again!! the latter one is pretty straightforward!!

OpenStudy (anonymous):

i get [1/2[-1/(2+3y) +1/3y]

OpenStudy (experimentx):

Oh ... yes ... i put the wrong limit, for x .. i guess that is right .. you can do the rest.

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

because when i continue i dont get the same eanswer

OpenStudy (experimentx):

try wolfram

OpenStudy (anonymous):

k let me simplify if then i will tell you what i get

OpenStudy (experimentx):

wolfram seems to comply with your answer unless you or I made mistake in integration ... you should the answer!

OpenStudy (anonymous):

i give up #sad